OpenStudy (anonymous):

Need help...Use the Laplace transform to solve differential equation: dx/dt + 7x = 5cos(2t) Initial conditions: x(0) = 4, x'(0) = -4 Assume forcing functions are 0 prior to t = 0-

OpenStudy (anonymous):

I did the first step taking the laplace of both sides. SX(s) - x(0-) + 7X(x) = 5s/(s^2 +4)

OpenStudy (anonymous):

x(0-) = 4 SX(s) - 4 + 7X(s) = 5s(s^2 + 4) (s + 7)X(s) = 4s^2 + 5s+ 16 / (s^2 +4) X(s) = (4s^2 + 5s + 16)/(s + 7)(s^2 + 4)

OpenStudy (anonymous):

So I guess my question is taking the inverse laplace transform and that has to do with partial fractions

OpenStudy (anonymous):

hey myininaya satellite helped me out with the partial fraction

OpenStudy (zarkon):

\[\frac{35s}{53(s^2+4)}+\frac{20}{53(s^2+4)}+\frac{177}{53(s+7)}\]

OpenStudy (anonymous):

thanks... I got the inverse laplace transform: 35/53cos(2t) + 10/53sin(2t) + 177/53e^(-7t)

OpenStudy (anonymous):

i wish i new jack about these transforms

OpenStudy (anonymous):

guess i need a book

OpenStudy (zarkon):

I don't teach this stuff enough to have all the transforms memorized...I always look them up

OpenStudy (anonymous):

yeah it really helps it turns differential equations into linear algebra makes it easy to solve. in this case I'm just using a table to substitue

OpenStudy (anonymous):

if it is like those stultifying formulas for "techniques of integration" or ode's i will forget about it

OpenStudy (anonymous):

but it seems to get asked fairly frequently so it must be pretty common

OpenStudy (anonymous):

so for instant laplace of (sin3t) = 3/(s^2 + 9)

OpenStudy (anonymous):

laplace of dx/dt = sX(s) - x(0)

OpenStudy (anonymous):

helps you solve ODE algebraically

OpenStudy (anonymous):

yeah someone tried to explain it to me. said it turns ode's into polynomials

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so like x' + 5x = t

OpenStudy (anonymous):

with initial conditions x(0) = 0 x'(0) = 0

OpenStudy (anonymous):

laplace [x'] = sX(s)) - x(0) laplace [5x] = 5X(s) laplace [t] = 1/s^2

OpenStudy (anonymous):

So you have: sX(s) - 0 + 5X(s) = 1/s^2

OpenStudy (anonymous):

(s+5)X(s) = 1/s^2 X(s) = 1/[(s^2)(s+5)]

OpenStudy (anonymous):

So to find the solution which is x(t) you take inverse laplace of the expression on the right which involves partial fractions.

OpenStudy (anonymous):

great, more partial fractions!

OpenStudy (anonymous):

1/5s^2 + 1/25(s+5) - 1/25s

OpenStudy (anonymous):

haha...yeah..So inverse laplace will give us x(t) = t/5 + 1/25e^(-5t) - 1/25u(t)

OpenStudy (anonymous):

apparently you get \[x(t)=\frac{1}{25}(5t=e^{-5t}-1)\]

OpenStudy (anonymous):

As you do more practice you'll get the hang of it.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[x(t)=\frac{1}{25}(5t+e^{-5t}-1)\]

OpenStudy (anonymous):

U(t) just represents the step function it is 1 when t > 0 and 0 when t <0

OpenStudy (anonymous):

is it fun or donkey work?

OpenStudy (anonymous):

lol it gets tiring when the laplace transforms get complicated with partial fractions...but that's when matlab comes in lol we are also learning to use that software.

OpenStudy (anonymous):

thanks for your help too

OpenStudy (dumbcow):

i think your initial conditions are off, it should be x(0) = 4, x'(0) = -23 http://www.wolframalpha.com/input/?i=solve+x%27+%2B+7x+%3D+5cos%282t%29%2C+x%280%29%3D4%2C+x%27%280%29%3D-23 i solved it but i didn't need to use LaPlace transform

OpenStudy (anonymous):

well you only need the first initial condition if you're solving with laplace because laplace of dx/dt only involves initial condition x(0) but i'm checking the work over again

OpenStudy (dumbcow):

oh ok, yeah thats true, the second condition isn't neccessary, i just noticed our solution x(t) didn't match up with x'(0) = -4

OpenStudy (anonymous):

Okay I went through solving using method of undetermined coefficients and for that the result for the particular solution is the same 35/53cos2t +10/53sin2t, and when I plug in initial condition of x(0 = 4 I get the same thing c1 is 177/53 so the other solution the homogenous is 177/53e^(-7t). in fact in both cases because we are dealing with a first oder DFQ the initial condition of x'(0) is not important.

OpenStudy (anonymous):

it's just giving in the problem.